Find the equation of a line with the properties given. Write the equation in the form indicated.
1.)Through (5, -5) and parallel to the line whose equation is y = -9x +7; slope-intercept form

Answer by lwsshak3(11628)

Find the equation of a line with the properties given. Write the equation in the form indicated.
1.)Through (5, -5) and parallel to the line whose equation is y = -9x +7; slope-intercept form.
***
slope-intercept form of equation: y=mx+b, m=slope, b=y-intercept
y=-9x+7
slope=-9
slope of line parallel to given line=-9
equation:
y=-9x+b
solve for b using coordinates of given points on the line, (5,-5)
-5=-9*5+b
b=40
equation:
y=-9x+40